OpenStudy (anonymous):

What is the 7th term of the geometric sequence where a1 = 1,024 and a4 = −16? 1 −0.25 −1 0.25

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@am!rah

OpenStudy (fanduekisses):

use\[an=a _{1}*r^{n-1}\]

OpenStudy (fanduekisses):

This is the formula for a geometric sequence, where a1= first term, r=common rate

OpenStudy (anonymous):

My problem is plugging them in :c

OpenStudy (fanduekisses):

do you know what the common rate is?

OpenStudy (fanduekisses):

opps it is actually called the "common ratio" my fault sorry

OpenStudy (anonymous):

Yeah, i do. i just dont know how i plug them in xD i get confused on which numbers to put where.

OpenStudy (fanduekisses):

ok so you are looking for the 7th term so n=7

OpenStudy (fanduekisses):

a1= the first term (you were given)

OpenStudy (anonymous):

Mhm. And i dunno r.

OpenStudy (fanduekisses):

r = the common ratio

OpenStudy (anonymous):

i think it might be -4.

OpenStudy (fanduekisses):

ok so plug in :)

OpenStudy (anonymous):

its giving me a really big number ;-; Im so confused.

OpenStudy (fanduekisses):

the common ratio might not be -4 then

OpenStudy (anonymous):

@satellite73

OpenStudy (jdoe0001):

\(\large { \begin{array}{rrllll} term&value \\\hline\\ a_1&1024\\ a_2&a_1\cdot r\\ a_3&a_2\cdot r\to (a_1\cdot r)\cdot r\\ a_4&a_3\cdot r\to (a_1\cdot r\cdot r)\cdot r\to -16 \end{array}\\ \quad \\\\ \quad \\ \implies a_1\cdot r^3=-16\implies r=? }\)

OpenStudy (anonymous):

What.

OpenStudy (jdoe0001):

\(\large { \begin{array}{llll} term&value \\\hline\\ a_1&1024\\ a_2&a_1\cdot r\\ a_3&a_2\cdot r\to (a_1\cdot r)\cdot r\\ a_4&a_3\cdot r\to (a_1\cdot r\cdot r)\cdot r\to -16 \end{array}\\ \quad \\\\ \quad \\ \implies a_1\cdot r^3=-16\implies r=? }\)

OpenStudy (jdoe0001):

"r" is the common multiplier, or "common ratio"

OpenStudy (anonymous):

Yeahh, i said -4.

OpenStudy (jdoe0001):

once you find the common ratio, and the 1st term, is given, then you can find pretty much any term in the sequence

OpenStudy (anonymous):

:c i still dont get how to plug it in. can you show me?

OpenStudy (jdoe0001):

well... is not -4.... recheck it

OpenStudy (anonymous):

-2?

OpenStudy (jdoe0001):

\(\bf a_1\cdot r^3=-16\implies 1024r^3=-16 \implies r=?\)

OpenStudy (anonymous):

I GOT -0.25 For the answer. ;-;

OpenStudy (jdoe0001):

well... how did you get -0.25 if you dunno the common ratio?

OpenStudy (anonymous):

i thought it was -4 D;

OpenStudy (jdoe0001):

well... let's try -4 r = -4 so that means \( \bf a_1\cdot (-4)^3=-16\implies 1024(-64)=-16\implies 65536\ne -16\)